o Determine shear stress for the bolt at A. - Diameter of the bolt is 15mm 400N 200 N K lm * 3m x
Answer #1

400N 200N À 3m calculating reactions VA + B = 400N. moment at A E MA = 0 -> - 400 x 4 + VB x3 = 0 - VB= 4 x 450 = 533:33 N AFree body diagram of point a - 200 N A Resultant force - - 133.33 N R=d133.33²+200² R = 240.37 Ny M - - - - - A Shearplane No